Wirh the following price and dividend information, calculate the sample variance for the returns. (Round to 4 decimals)

Year

Price

Dividend

0

49.72

1

53.54

1.00

2

59.22

1.10

3

51.3

1.20

4

53.45

1.30

5

56.35

1.40
Answer #1
Variance 0.9003%
Year Price Dividend Return=(Dividend+P1-P0)/P0 (X-Average)^2
0 49.72
1 53.54 1 9.69% 0.00206
2 59.22 1.1 12.66% 0.005636
3 51.3 1.2 -11.35% 0.027237
4 53.45 1.3 6.73% 0.000246
5 56.35 1.4 8.04% 0.000835
Average 5.1560%
Total 0.036013

Variance = Total (X-Average)^2 / (n-1) = 0.036013/(5-1) =

0.9003%